OpenStudy (anonymous):

Solve. 2x^2 + x = 15.

OpenStudy (anonymous):

ANSWER CHOICES x = 5/2 and x = -3 x = 2 and x = -15 x = -2 and x = 5/3 x = 17 and x = 13

OpenStudy (anonymous):

2x^2 - x + 15=(2x+5)(x-3)

OpenStudy (anonymous):

so A ?

OpenStudy (nurali):

2x^2 + x = 15 2x^2 +x -15=0 2x^2 +6x-5x-15=0 2x(x+3)-5(x+3)=0 (2x-5)(x+3)=0 2x-5=0 or x+3=0 x=5/2 or x=-3